# 3D Feature Distillation with Object-Centric Priors

### v4

@@ A.7 Robot Experiments
Implementation We develop our language-guided grasping behavior in ROS, using DROP-CLIP for grounding the user’s query and RGB-D grasp detection network, HGGD [43], for generating 6-DoF grasp proposals. Our pipeline is shown in Fig. 17. We process the raw sensor point-cloud with RANSAC from open3d library with distance threshold $0.1$, ransac_n=3 and 1000 iterations to segment out the table points, and then upscale to ×$\times 10$. We use in-scene category names as negative prompts and do inference with a threshold of $0.7$. To match the grounded object points with grasp proposals, we transform predicted grasps to world frame and move their center at the gripper’s tip. We then calculate euclidean distances between the gripper’s tip and the thresholded prediction’s center. In real robot experiments, we run statistical outlier removal from open3d with neighbor_size=25 and std_ratio=2.0 to remove noisy points from the prediction’s center. The top-3 closest grasps are given as goal for an inverse kinematics motion planner. We manually mark grasp attempts as success/failure in real robot and leverage the simulator state to do it automatically in Gazebo. Visualizations of simulated / real robot trials are illustrated in Fig. 18, while related videos are included as supplemetary material.
@@ A.8 Detailed Related Work
3D CLIP Feature Distillation OpenScene [12] distills OpenSeg [9] multi-view features with a point-cloud encoder, while follow-up work Open3DSG [15] extends to scene graph generation by further distilling object-pair representations from other vision-language foundation models [83] as graph edges. CLIP-FO3D [21] replaces OpenSeg pixel-wise with multi-scale crops from CLIP to further enhance generalization. All above works use dense 2D features and fuse point-wise, thus suffering from ‘patchyness’ issue. Further, these works focus on 3D room scan data, and hence do not have to deal with the effects of clutter in the multi-view fusion process. LERF [13] replaces point-cloud encoders with neural fields, and distils multi-scale crop CLIP features into a continuous feature field that can provide features in any region of the input space. The authors deal with the ‘patchyness’ issue using DINO regularization. Similar works OpenNeRF [19] and F3RM [16] use MaskCLIP to extract features and avoid DINO-regularization. All above works make the assumption that all views are equally informative and rely on dense number of views at test-time to resolve the noise in the distilled features. LangSplat [24] is a similar Gaussian splatting work that exploits spatial priors (i.e. segmentation masks) to distill object-level CLIP features but does not perform view selection based on semantics, and lies in the same general family of works as fields, i.e., requiring multiple camera images and computational resources to work at test-time. In contrast, our method reconstructs semantically-informed, view-independent 3D features from single-view RGB-D inputs, while enabling real-time inference.

### v5

@@ A.7 Robot Experiments
Figure 20: Visualization of grounding queries in real robot trials, for baseline method MaskCLIP→3D (bottom) and our DROP-CLIP (top), where we ensemble the predictions of our method with the 2D baseline. DROP-CLIP produces more robust features (middle column) which lead to crispier segmentation (right column.)
Implementation We develop our language-guided grasping behavior in ROS, using DROP-CLIP for grounding the user’s query and RGB-D grasp detection network, HGGD [44], for generating 6-DoF grasp proposals. Our pipeline is shown in Fig. 18. We process the raw sensor point-cloud with RANSAC from open3d library with distance threshold $0.1$, ransac_n=3 and 1000 iterations to segment out the table points, and then upscale to $\times 10$. We use in-scene category names as negative prompts and do inference with a threshold of $0.7$. To match the grounded object points with grasp proposals, we transform predicted grasps to world frame and move their center at the gripper’s tip. We then calculate euclidean distances between the gripper’s tip and the thresholded prediction’s center. In real robot experiments, we run statistical outlier removal from open3d with neighbor_size=25 and std_ratio=2.0 to remove noisy points from the prediction’s center. The top-3 closest grasps are given as goal for an inverse kinematics motion planner. We manually mark grasp attempts as success/failure in real robot and leverage the simulator state to do it automatically in Gazebo. Visualizations of simulated / real robot trials are illustrated in Fig. 19, experiments with grounding different objects with fine-grained attributes in Fig. 20, while related videos are included as supplemetary material.
@@ A.8 Detailed Related Work
Offline 3D CLIP Feature Distillation OpenScene [12] distills OpenSeg [9] multi-view features with a point-cloud encoder, while follow-up work Open3DSG [15] extends to scene graph generation by further distilling object-pair representations from other vision-language foundation models [87] as graph edges. CLIP-FO3D [21] replaces OpenSeg pixel-wise features with multi-scale crops from CLIP to further enhance generalization. All above works use dense 2D features and fuse point-wise, thus suffering from ‘patchyness’ issue. Further, these works distill features using 3D room scan data [22, 1], which lack diverse object catalogs and do not have to deal with the effects of clutter in the multi-view fusion process, as we do with the introduction of MV-TOD.
Online 3D CLIP Feature Distillation LERF [13] replaces point-cloud encoders with neural fields, and distils multi-scale crop CLIP features into a continuous feature field that can provide features in any region of the input space. The authors deal with the ‘patchyness’ issue using DINO regularization. Similar works OpenNeRF [19] and F3RM [16] use MaskCLIP to extract features and avoid DINO-regularization. All above works make the assumption that all views are equally informative and rely on dense number of views at test-time to resolve the noise in the distilled features. A more recent line of works replace NeRFs with 3D Gaussian Splatting (3DGS) [88] to improve inference time and memory consumption and perform similar feature distillation from LSeg, OpenSeg or CLIP multi-scale crops. Similar to our work, some 3DGS approaches [24, 75, 76, 77] also exploit spatial priors (i.e. segmentation masks) to distill object-level CLIP features, but do not perform view selection based on semantics. Further, 3DGS approaches lie in the same general family of works as fields, i.e., online distillation in specific scenes, requiring multiple camera images and computational resources to work at test-time. In contrast, our method is distilled offline in MV-TOD to reconstruct semantically-informed, view-independent 3D features from single-view RGB-D inputs, can be applied zero-shot in novel scenes without training, and enables real-time inference.